# Universal Hitchin moduli spaces

### v1

@@ 5.4. Metric structure of the universal moduli of Hitchin’s equations
In this section we combine the results of the previous two sections in order to define a universal moduli spaces of solutions of Hitchin’s equations (2.15), varying over (a cover of) Teichmuller space. As in the previous sections, we consider a fixed principal $K$-bundle $E_{K}$ over a smooth compact oriented surface $\Sigma$ with fixed symplectic form $\omega$.
@@
We say that a triple $(J,A,\psi)\in\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$ is a solution of the coupled Hitchin’s equations, if the following conditions are satisfied
@@
By the explicit formula for the moment in Proposition 5.10, the underlying space to $\mathcal{U}^{Hit}(G)_{\alpha}^{\varepsilon}$ is independent of the choice of parameters, as the coupling terms in the scalar equation in (5.15) vanishes identically upon imposing the integrability condition $\bar{\partial}_{J,A}\varphi=0$. We shall denote this moduli space simply by $\mathcal{U}^{Hit}(G)$. Hence, for any $\alpha>0$ and $\varepsilon\in\{-1,1\}$, the moduli space $\mathcal{U}^{Hit}(G)$ parametrizes solutions of the equations (5.16) modulo the $\widetilde{{\curly K}}-$action, and the only difference is the induced presymplectic structure on the moduli space. This structure is furthermore symplectic in the case $\varepsilon=-1$, by Lemma 5.8. By construction, there are natural maps (cf. (5.3))
$\mathcal{U}^{Hit}(G)\to\mathcal{U}^{Higgs}(G)\to\mathcal{T}.$
Hence, in particular, $\mathcal{U}^{Hit}(G)$ can be regarded as a fibration over the Teichmüller space $\mathcal{T}$. Similarly as in Section 4.2, since the symmetric $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is not positive definite (see Lemma 5.8), it is not obvious a priori that $\mathcal{U}^{Hit}(G)$ inherits a complex structure compatible with the (pre)symplectic structure induced by $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$. The main goal of this section is to study sufficient conditions under which this natural condition for the moduli space holds, furthermore proving that the map $\mathcal{U}^{Hit}(G)\to\mathcal{U}^{Higgs}(G)$ is holomorphic.
The analysis is analogue to that in Sections 4.2 and 3.3, and hence we just outline the argument. The first step is to undertake a gauge fixing for solutions of the coupled Hitchin’s equations (5.16), whereby the complex structure (5.1) and the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ descend to the moduli space. We observe that this will produce a priori different complex structures on $\mathcal{U}^{Hit}(G)$.
We fix a solution $(J,A,\psi)$ of the coupled Hitchin’s equations (5.16).
@@
The linearization of the coupled Hitchin’s equations (5.16) at $(J,A,\psi)$ is given by
@@
(5.18) ${(\mathcal{S}^{*}_{0})}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{S}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{P}}$${\mathcal{S}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\bf L}}$${\mathcal{S}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
The cohomology $H^{1}(\mathcal{S}^{*}_{0}):=\frac{\operatorname{Ker}{\bf L}}{\operatorname{Im}\,\mathbf{P}}$ can be formally identified with the tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{Hit}(G)$. Our next result shows that the moduli space $\mathcal{U}^{Hit}(G)$ is finite dimensional.
@@
In order to construct a complex structure induced by (5.1) on the moduli space, we work orthogonal to the image of the infinitesimal action operator $\mathbf{P}$ in (5.18) with respect to the indefinite pairing $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$. The existence of this complex structure will automatically yield a symmetric tensor of type $(1,1)$, since the two-form $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$ is well defined on the cohomology $H^{1}(\mathcal{S}^{*}_{0})$ by Proposition 5.10. Consider the $L^{2}$-pairing on $\mathcal{S}^{0}$ define by (4.8). Consider the map $\mu\colon\curly{X}\to\mathcal{S}^{0}$ defined by (cf. Proposition 5.10)
$\mu^{\mathbb{I}}(J,A,\psi)=\left(\mu^{\mathbb{I}}_{0},\mu^{\mathbb{I}}_{1}\right).$
@@
$\mu^{\mathbb{I}}_{0}$ $=-\left(\varepsilon S_{J}\omega+\alpha d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-\alpha d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right)$
@@
The following operator provides a formal adjoint of the infinitesimal action $\mathbf{P}$ for the pairings (4.8) and (5.11)
$\mathbf{P}^{*}_{\alpha,\varepsilon}=\delta\mu^{\mathbb{I}}\circ\mathbb{I}\colon\mathcal{S}^{1}\to\mathcal{S}^{0}.$
@@
The key condition on the solution $(J,A,\psi)$ of (5.16) which we need to assume in order to construct the complex structure on the moduli space is the vanishing of the kernel of $\mathcal{L}_{0}^{\alpha,\varepsilon}$. Notice that, unlike in the standard cases in gauge theory in which the parameter space metric is positive definite, $\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}$ does not relate in general to automorphisms of the triple $(J,\psi,A)$, but rather to null vectors with respect to $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$.
@@
The operator $\mathcal{L}_{0}^{\alpha,\varepsilon}$ is Fredholm with zero index. Furthermore, elements $(f,u)\in\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}$ are smooth.
Assuming that $\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}$ is trivial, in the next result we obtain a natural gauge fixing via a $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$-orthogonal decomposition
@@
Assume that $\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}$. Then, there exists an orthogonal decomposition (5.20) for the pairing $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$. Consequently, for any element $v\in\mathcal{S}^{1}$ there exists a unique $\Pi v\in\operatorname{Im}\;\mathbf{P}$ such that $(\dot{J},a,\dot{\psi})=v-\Pi v$ solves the linear equation
@@
$\mathcal{H}^{1}(\mathcal{S}^{*}_{0})=\ker\bf{L}\cap\ker{\bf L}\circ\mathbb{I}.$
Our next result provides our gauge fixing mechanism for the linearization of the coupled Hitchin’s equations (5.16).
@@
Assume $\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}$. Then, the inclusion $\mathcal{H}^{1}(\mathcal{S}^{*}_{0})\subset\ker\bf{L}$ induces an isomorphism
$\mathcal{H}^{1}(\mathcal{S}^{*}_{0})\simeq H^{1}(\mathcal{S}^{*}_{0}).$
@@
$\mathcal{U}^{*}_{\alpha,\varepsilon}=\{[(J,A,\psi)]\;|\;\operatorname{ker}\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}\}\subset\mathcal{U}^{Hit}(G).$
@@
(5.22) $\begin{split}\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B(a_{1}\wedge a_{2})-\alpha\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})\\ &-\frac{\alpha}{2}\int_{\Sigma}B\left(\psi(J\dot{J}_{1})\wedge\dot{\psi}_{2}\right)-\frac{\alpha}{2}\int_{\Sigma}B\left(\dot{\psi}_{1}\wedge\psi(J\dot{J}_{2})\right)\\ &-\frac{\alpha}{2}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right),\\ \mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B(a\wedge Ja)+\alpha\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{\alpha}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}_{2}+\tfrac{\alpha}{2}\psi(J\dot{J}_{2})\right)\right)\\ &-\frac{\alpha}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right),\end{split}$
@@ 5.5. Comparison with $\boldsymbol{\mathcal{U}^{Higgs}(G)}$ and existence
$\mathcal{U}^{Hit}(G)\supset\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{Higgs}(G)$
and hence a holomorphic map into Teichmuller space $\mathcal{T}$. We will also prove that for genus of the surface $\Sigma$ bigger than zero, the open $\mathcal{U}^{*}\subset\mathcal{U}^{Hit}(G)$ is non-empty for sufficiently small values of $\alpha$. Our analysis is very similar to that in Section 4.3, and hence we omit the details.
@@
(5.23) $\begin{split}\bar{\partial}_{J,A}\psi^{1,0_{J}}=0\end{split}$
@@
(5.24) ${(\mathcal{C}_{0}^{*})}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{C}^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{P}^{c}_{0}}$${\mathcal{C}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\bf L^{c}_{0}}}$${\mathcal{C}^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
@@
$\mathcal{C}^{0}=\operatorname{Lie}\widetilde{\curly{G}}\cong\Omega^{0}(T\Sigma)\oplus\Omega^{0}(X,E_{G}(\mathfrak{g})),\qquad\mathcal{C}^{1}=\mathcal{S}^{1},\qquad\mathcal{C}^{2}=\Omega^{2}(X,E_{G}(\mathfrak{g}))$
@@
Applying Theorem 2.7, a solution $(J,A,\psi)$ of the coupled Hitchin’s equations (5.16) induces a polystable $G$-Higgs bundle $(E,\varphi)\cong(\bar{\partial}_{J,A},\psi^{1,0_{J}})$. This fact, jointly with the natural inclusion $\widetilde{{\curly K}}\subset\widetilde{\curly{G}}$, leads to a continuous map
(5.25) $\mathcal{U}^{Hit}(G)\to\mathcal{U}^{Higgs}(G)\colon[(J,A,\psi)]\mapsto[(J,A,\psi)].$
@@
(5.26) ${\mathcal{U}^{Hit}(G)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{U}^{Higgs}(G)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mu_{\curly{H}}^{-1}(0)/\curly{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}}$
where $\mu_{\curly{H}}^{-1}(0)/\curly{H}$ is the moduli space of constant scalar curvature Kähler metrics with total volume $V$, modulo $\omega$-Hamiltonian diffeomorphisms. The fibre of $\mu_{\curly{H}}^{-1}(0)/\curly{H}\to\mathcal{T}$ can be identified with $H^{1}(\Sigma,\mathbb{R})$ [F]. Building on Theorem 5.18, our next goal is to prove that this induces a holomorphic map $\mathcal{U}^{Hit}(G)\supset\mathcal{U}^{*}\to\mathcal{U}^{Higgs}(G)$.
@@
(5.27) ${H^{1}(\mathcal{S}^{*}_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\mathcal{C}^{*}_{0}).}$
@@
To finish this section, we address the question of non-emptyness of the moduli space $\mathcal{U}^{Hit}(G)$, in genus $g(\Sigma)\geqslant 2$. For this, we fix a compact Riemann surface $X=(\Sigma,J)$ and consider a $G$-Higgs bundle $(E,\varphi)$. In this setup, consider the coupled equations
(5.28) $\begin{split}F_{h}-[\varphi\wedge\tau_{h}\varphi]&=0\\ \varepsilon S_{g}&=\varepsilon\frac{2\pi\chi(\Sigma)}{V}\end{split}$
for pairs $(g,h)$, where $g$ is a Kähler metric on $X$ with total volume $V$ and $h\in\Omega^{0}(E_{G}(G/K))$ is a reduction of structure group of $E$ to $K$. A solution $(g,h)$ of the equations (5.28) determines then a solution of (5.16), given by the triple $(J,A_{h},-i(\varphi-\tau_{h}\varphi))$. We are ready to present our last main result.
@@
(5.29) $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mu_{\curly{H}}^{-1}(0)/\curly{H}$
@@
The restriction of the $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ in Corollary 5.19 to the fibres of (5.29) is $\alpha\mathbf{g}$, where $\mathbf{g}$ denotes the hyperkähler metric on the moduli space of solutions of Hitchin’s equations. Consequently, (5.29) has a natural structure of Kähler fibration with coupling form $\boldsymbol{\omega}_{1,0}$ (see (5.22)) and Kähler Ehresmann connection.
@@
The proof is analogue of that of Theorem 4.14, and is therefore omitted. The only difference is that we can control the non-degeneracy of $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ along the fibres over the moduli space of constant scalar curvature Kähler metrics $\mathcal{N}=\mu_{\curly{H}}^{-1}(0)/\curly{H}$. To see this, we observe that if $(\dot{J},a,\dot{\psi})$ is in the tangent to the fibre over $\mathcal{N}$, then due to the gauge fixing in Proposition 5.17, we must have $\dot{J}=0$, and the statement follows from Corollary 5.19. Hence, assuming that $\dot{J}=L_{J\eta_{f}}J$, and since the kernel of the Lichnerowicz operator is given by the Hamiltonian functions of Killing holomorphic vector field, $f$ must be constant by our assumption $g(\Sigma)\geqslant 2$. ∎

### v2

@@ 5.4. Metric structure of the universal moduli of Hitchin’s equations
In this section we combine the results of the previous two sections to define a universal moduli space of solutions of Hitchin’s equations (2.12), varying over (a cover of) the Teichmüller space. As in the previous sections, we consider a fixed principal $K$-bundle $E_{K}$ over a smooth compact oriented surface $\Sigma$ with fixed symplectic form $\omega$.
@@
A triple $(J,A,\psi)\in\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$ is a solution of the coupled Hitchin equations, if the following conditions are satisfied
@@
By the explicit formula for the moment map in Proposition 5.10, the underlying space of $\mathcal{U}^{Hit}(G)_{\alpha}^{\varepsilon}$ is independent of the choice of parameters, since the ‘coupling term’ in the scalar equation in (5.15) vanishes identically when the integrability condition $\bar{\partial}_{J,A}\varphi=0$ is imposed. We shall denote this moduli space simply by $\mathcal{U}^{Hit}(G)$. Hence, for any $\alpha>0$ and $\varepsilon\in\{-1,1\}$, the moduli space $\mathcal{U}^{Hit}(G)$ parametrizes solutions of the equations (5.16) modulo the $\widetilde{{\curly K}}-$action, and the only difference is the induced presymplectic structure on the moduli space. This structure is furthermore symplectic in the case $\varepsilon=-1$, by Lemma 5.8. By construction, there are natural maps (cf. (5.3))
$\mathcal{U}^{Hit}(G)\longrightarrow\mathcal{U}^{Higgs}(G)\longrightarrow\mathcal{T}.$
Hence, in particular, $\mathcal{U}^{Hit}(G)$ can be regarded as a fibration over the Teichmüller space $\mathcal{T}$. As in Section 4.2, since the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is not positive definite (see Lemma 5.8), it is not obvious a priori whether $\mathcal{U}^{Hit}(G)$ inherits a complex structure compatible with the (pre)symplectic structure induced by $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$. The main goal of this section is to study sufficient conditions under which this natural condition holds for the moduli space, and to prove that the map $\mathcal{U}^{Hit}(G)\to\mathcal{U}^{Higgs}(G)$ is holomorphic.
The analysis is analogous to that in Sections 4.2 and 3.3, so we only outline the argument. The first step is to undertake a gauge fixing for solutions of the coupled Hitchin equations (5.16), whereby the complex structure (5.1) and the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ descend to the moduli space. We observe that this will produce a priori different complex structures on $\mathcal{U}^{Hit}(G)$.
We fix a solution $(J,A,\psi)$ of the coupled Hitchin equations (5.16).
@@
The linearization of the coupled Hitchin equations (5.16) at $(J,A,\psi)$ is given by
@@
[…]
The cohomology $H^{1}(\mathcal{S}^{*}_{0}):=\frac{\ker{\bf L}}{\operatorname{Im}\,\mathbf{P}}$ can be formally identified with the tangent space $T_{[(J,A,\psi)]}\mathcal{U}^{Hit}(G)$. Our next result shows that the moduli space $\mathcal{U}^{Hit}(G)$ is finite dimensional.
@@
In order to construct a complex structure induced by (5.1) on the moduli space, we work orthogonal to the image of the infinitesimal action operator $\mathbf{P}$ in (5.18) with respect to the indefinite pairing $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$. The existence of this complex structure will automatically yield a symmetric tensor of type $(1,1)$, since the 2-form $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$ is well defined on the cohomology $H^{1}(\mathcal{S}^{*}_{0})$ by Proposition 5.10. Consider the $L^{2}$-pairing on $\mathcal{S}^{0}$ define by (4.8). Consider the map $\mu\colon\curly{X}\to\mathcal{S}^{0}$ defined by (cf. Proposition 5.10)
$\mu^{\mathbb{I}}(J,A,\psi)=\left(\mu^{\mathbb{I}}_{0},\mu^{\mathbb{I}}_{1}\right),$
@@
$\mu^{\mathbb{I}}_{0}$ $=-\left(\varepsilon S_{J}\omega+\alpha d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-\alpha d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right),$
@@
The following operator provides a formal adjoint of the infinitesimal action $\mathbf{P}$ for the pairings (4.8) and (5.11):
$\mathbf{P}^{*}_{\alpha,\varepsilon}=\delta\mu^{\mathbb{I}}\circ\mathbb{I}\colon\mathcal{S}^{1}\longrightarrow\mathcal{S}^{0}.$
@@
The key condition on the solution $(J,A,\psi)$ of (5.16) which we need to assume in order to construct the complex structure on the moduli space is the vanishing of the kernel of $\mathcal{L}_{0}^{\alpha,\varepsilon}$. Notice that, unlike in the standard cases in gauge theory in which the parameter space metric is positive definite, $\ker\mathcal{L}_{0}^{\alpha,\varepsilon}$ does not relate in general to automorphisms of the triple $(J,\psi,A)$, but rather to null vectors with respect to $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$.
@@
The operator $\mathcal{L}_{0}^{\alpha,\varepsilon}$ is Fredholm with zero index. Furthermore, elements $(f,u)\in\ker\mathcal{L}_{0}^{\alpha,\varepsilon}$ are smooth.
Assuming that $\ker\mathcal{L}_{0}^{\alpha,\varepsilon}$ is trivial, in the next result we obtain a natural gauge fixing via a $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$-orthogonal decomposition
@@
Assume that $\ker\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}$. Then, there exists an orthogonal decomposition (5.20) for the pairing $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$. Consequently, for any element $v\in\mathcal{S}^{1}$ there exists a unique $\Pi v\in\operatorname{Im}\;\mathbf{P}$ such that $(\dot{J},a,\dot{\psi})=v-\Pi v$ solves the linear equation
@@
$\mathcal{H}^{1}(\mathcal{S}^{*}_{0})=\ker\bf{L}\cap\ker({\bf L}\circ\mathbb{I}).$
Our next result provides our gauge fixing mechanism for the linearization of the coupled Hitchin equations (5.16).
@@
Assume $\ker\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}$. Then, the inclusion $\mathcal{H}^{1}(\mathcal{S}^{*}_{0})\subset\ker\bf{L}$ induces an isomorphism
$\mathcal{H}^{1}(\mathcal{S}^{*}_{0})\cong H^{1}(\mathcal{S}^{*}_{0}).$
@@
$\mathcal{U}^{*}_{\alpha,\varepsilon}=\{[(J,A,\psi)]\;|\;\ker\mathcal{L}_{0}^{\alpha,\varepsilon}=\{0\}\}\subset\mathcal{U}^{Hit}(G).$
@@
(5.22) $\begin{split}\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B(a_{1}\wedge a_{2})-\alpha\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})\\ &-\frac{\alpha}{2}\int_{\Sigma}B\left(\psi(J\dot{J}_{1})\wedge\dot{\psi}_{2}\right)-\frac{\alpha}{2}\int_{\Sigma}B\left(\dot{\psi}_{1}\wedge\psi(J\dot{J}_{2})\right)\\ &-\frac{\alpha}{2}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right),\\ \mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B(a\wedge Ja)+\alpha\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{\alpha}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}_{2}+\tfrac{\alpha}{2}\psi(J\dot{J}_{2})\right)\right)\\ &-\frac{\alpha}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right).\end{split}$
@@ 5.5. Comparison with $\boldsymbol{\mathcal{U}^{Higgs}(G)}$ and existence
$\mathcal{U}^{Hit}(G)\supset\mathcal{U}^{*}_{\alpha,\varepsilon}\longrightarrow\mathcal{U}^{Higgs}(G),$
and hence a holomorphic map into Teichmüller space $\mathcal{T}$. We will also prove that for genus of the surface $\Sigma$ bigger than zero, the open $\mathcal{U}^{*}\subset\mathcal{U}^{Hit}(G)$ is non-empty for sufficiently small values of $\alpha$. Our analysis is very similar to that in Section 4.3, and hence we omit the details.
@@
(5.23) $\begin{split}\bar{\partial}_{J,A}\psi^{1,0_{J}}=0.\end{split}$
@@
[…]
@@
$\mathcal{C}^{0}=\operatorname{Lie}\widetilde{\curly{G}}\cong\Omega^{0}(T\Sigma)\oplus\Omega^{0}(\Sigma,E_{G}(\mathfrak{g})),\qquad\mathcal{C}^{1}=\mathcal{S}^{1},\qquad\mathcal{C}^{2}=\Omega^{2}(\Sigma,E_{G}(\mathfrak{g}))$
@@
Applying Theorem 2.7, a solution $(J,A,\psi)$ of the coupled Hitchin equations (5.16) induces a polystable $G$-Higgs bundle $(E,\varphi)\cong(\bar{\partial}_{J,A},\psi^{1,0_{J}})$. This fact, jointly with the natural inclusion $\widetilde{{\curly K}}\subset\widetilde{\curly{G}}$, leads to a continuous map
(5.25) $\begin{array}[]{rcl}\mathcal{U}^{Hit}(G)&\longrightarrow&\mathcal{U}^{Higgs}(G)\\ \left[(J,A,\psi)\right]&\longmapsto&\left[(J,A,\psi)\right].\end{array}$
@@
(5.26) $\mathcal{U}^{Hit}(G)$$\mathcal{U}^{Higgs}(G)$$\mu_{\curly{H}}^{-1}(0)/\curly{H}$$\mathcal{T}$}
where $\mu_{\curly{H}}^{-1}(0)/\curly{H}$ is the moduli space of constant scalar curvature Kähler metrics with total volume $V$, modulo $\omega$-Hamiltonian diffeomorphisms. The fibre of $\mu_{\curly{H}}^{-1}(0)/\curly{H}\to\mathcal{T}$ can be identified with $H^{1}(\Sigma,\mathbb{R})$ [Fu]. Building on Theorem 5.18, our next goal is to prove that this induces a holomorphic map $\mathcal{U}^{Hit}(G)\supset\mathcal{U}^{*}\to\mathcal{U}^{Higgs}(G)$.
@@
(5.27) $H^{1}(\mathcal{S}^{*}_{0})$$H^{1}(\mathcal{C}^{*}_{0}),$}
@@
To finish this section, we address the question of non-emptyness of the moduli space $\mathcal{U}^{Hit}(G)$, in genus $g(\Sigma)\geqslant 2$. For this, we fix a compact Riemann surface $X=(\Sigma,J)$ and consider a $G$-Higgs bundle $(E,\varphi)$. In this setup, we consider the coupled equations
(5.28) $\begin{split}F_{h}-[\varphi\wedge\tau_{h}\varphi]&=0,\\ \varepsilon S_{g}&=\varepsilon\frac{2\pi\chi(\Sigma)}{V},\end{split}$
for pairs $(g,h)$ consisting of a Kähler metric $g$ on $X$ with total volume $V$ and a reduction $h\in\Omega^{0}(E_{G}(G/K))$ of the structure group of $E$ to $K$. A solution $(g,h)$ of the equations (5.28) determines then a solution of (5.16), given by the triple $(J,A_{h},-i(\varphi-\tau_{h}\varphi))$. We are ready to present our last main result.
@@
(5.29) $\mathcal{U}^{*}_{\alpha,\varepsilon}\longrightarrow\mu_{\curly{H}}^{-1}(0)/\curly{H}$
@@
The restriction of the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ in Corollary 5.19 to the fibres of (5.29) is $\alpha\mathbf{g}$, where $\mathbf{g}$ denotes the hyperkähler metric on the moduli space of solutions of Hitchin’s equations. Consequently, (5.29) has a natural structure of Kähler fibration with coupling form $\boldsymbol{\omega}_{1,0}$ (see (5.22)) and Kähler Ehresmann connection.
@@
The proof is analogous to that of Theorem 4.14 and is therefore omitted. The only difference is that we can control the non-degeneracy of $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ along the fibres over the moduli space of constant scalar curvature Kähler metrics $\mathcal{N}=\mu_{\curly{H}}^{-1}(0)/\curly{H}$. To see this, we observe that if $(\dot{J},a,\dot{\psi})$ is in the tangent to the fibre over $\mathcal{N}$, then due to the gauge fixing in Proposition 5.17, we must have $\dot{J}=0$, and the statement follows from Corollary 5.19. Hence, assuming that $\dot{J}=L_{J\eta_{f}}J$, and since the kernel of the Lichnerowicz operator is given by the Hamiltonian functions of Killing holomorphic vector field, $f$ must be constant by our assumption $g(\Sigma)\geqslant 2$. ∎